What is the molar solubility of silver chloride in pure water if its Ksp=1.8*10^-10
Answer #1

At equilibrium:

AgCl <----> Ag+ + Cl-

   s s

Ksp = [Ag+][Cl-]

1.8*10^-10=(s)*(s)

1.8*10^-10= 1(s)^2

s = 1.342*10^-5 M

Answer: 1.3*10^-5 M